# Designing an Optimal Sensor Network via Minimizing Information Loss

### v1

@@ 3.2 Connection to Sparse Variational Gaussian Processes
Our first assumption regards signal-to-noise ratios, allowing us to replace the inducing point means $\mu=\mathbb{E}[q(\boldsymbol{\mathbf{u}})]$ with observations. Recall the closed-form solution of $\mu$,
$\sigma^{-2}\boldsymbol{\mathbf{K}}_{MM}\mathopen{}\mathclose{{\left(\boldsymbol{\mathbf{K}}_{MM}+\sigma^{-2}\boldsymbol{\mathbf{K}}_{MN}\boldsymbol{\mathbf{K}}_{NM}}}\right)^{-1}\boldsymbol{\mathbf{K}}_{MN}\boldsymbol{\mathbf{y}},$ (21)
where $\sigma^{2}$ is determined as the estimated noise variance plus a regularizer $\operatorname{tr}{\boldsymbol{\mathbf{K}}_{\text{post}}}/N.$ So long as the sparse variational approximation is good, we expect $\operatorname{tr}{\boldsymbol{\mathbf{K}}_{\text{post}}}/N$ to be small, and thus $\sigma^{2}$ to be close to the actual estimated noise level. Then, with a high signal-to-noise ratio in the simulation, an observation $y(\boldsymbol{\mathbf{z}}_{m})$ will approximately be $\boldsymbol{\mathbf{\mu}}_{m}$.
If we additionally assume a high signal-to-noise ratio in observations of the real process, we can thus approximate the corresponding $\boldsymbol{\mathbf{\mu}}_{m}$. This is a surprisingly realistic assumption in many settings: for example, in temperature observations, noise may be on the scale of a few tenths of a degree Celsius.
@@
We proceed with a perturbative analysis, investigating the effect of data perturbation
$\delta_{\boldsymbol{\mathbf{Y}}}\triangleq\tilde{\boldsymbol{\mathbf{Y}}}_{N}-\boldsymbol{\mathbf{Y}}_{N}.$ (22)
In the collapsed bound Eq. 17, the data enters only through the “data fit” term
$D(\mathbf{Y}_{N})\triangleq\boldsymbol{\mathbf{Y}}_{N}^{\top}(\boldsymbol{\mathbf{Q}}_{NN}+\sigma^{2}_{\text{obs}}\mathbf{I})^{-1}\boldsymbol{\mathbf{Y}}_{N}.$ (23)
@@
through perturbations of the data fit term. This is clearly a quadratic form, which gives that
$\Delta D=\mathopen{}\mathclose{{\left[\nabla D(\boldsymbol{\mathbf{Y}}_{N})}}\right]\delta_{\boldsymbol{\mathbf{Y}}}+D(\delta_{\boldsymbol{\mathbf{Y}}}),$ (25)
where $\nabla D(\boldsymbol{\mathbf{Y}}_{N})=2(\boldsymbol{\mathbf{Q}}_{NN}+\sigma^{2}_{\text{obs}})^{-1}\boldsymbol{\mathbf{Y}}_{N}$.
@@
    A linear term, $\mathopen{}\mathclose{{\left[\nabla D(\boldsymbol{\mathbf{Y}}_{N})}}\right]\delta_{\boldsymbol{\mathbf{Y}}}$. There are two notable ways for this term to be small. The first is that both $\nabla D(\boldsymbol{\mathbf{Y}}_{N})$ and $\delta_{\boldsymbol{\mathbf{Y}}}$ are small, which can occur when the simulator has a smooth structure and is accurate (hence $\delta_{\boldsymbol{\mathbf{Y}}}$ is small). Potentially more interesting is when $\nabla D(\boldsymbol{\mathbf{Y}}_{N})$ and $\delta_{\boldsymbol{\mathbf{Y}}}$ have orthogonal structures, which may be interpreted as the real data varying in ways that are well-captured by the covariance function.
@@
    A quadratic term, $D(\delta_{\boldsymbol{\mathbf{Y}}})$. This term is small precisely when $\delta_{\boldsymbol{\mathbf{Y}}}$ is small, with a correlation similar to $(\boldsymbol{\mathbf{Q}}_{NN}+\sigma^{2}_{\text{obs}}\mathbf{I})$.
In total, this indicates that the simulator error is structured in such a way that it generally agrees with the covariance $(\boldsymbol{\mathbf{Q}}_{NN}+\sigma^{2}_{\text{obs}}\mathbf{I})$. Unfortunately, it is difficult to test this directly, but it can be informed via domain expertise or empirical evaluations. This limitation is inherent to our “pre-observation” setting, as no reliable statistical test can be designed without data we learn from subsequent deployment. Nonetheless, we provide the following heuristics to help understand when this may be the case:
@@
    High signal-to-noise ratio: The outputs of the simulator should dominate internal numerical noise or model uncertainty. This can be checked empirically by computing the signal-to-noise ratio across simulated realizations or by estimating it during GP training.
@@
    Non-localized errors: The simulator may exhibit errors, but those should be spatially diffuse rather than concentrated in specific regions. Our perturbative analysis (Section 3.2) shows that non-localized errors tend to have little effect on the ELBO and the resulting inducing points.
@@
    Coherence with sparse observations: If a limited number of real-world measurements are available, comparing them to simulator outputs can help establish empirical plausibility. In particular, modest correlation or bounded deviations can support the assumption that simulation-informed placement remains meaningful.
Note that small and non-localized errors may not be sufficient for a simulator to be useful in real-world decision making without the collection of real data. One potential reason for this is computational constraints: simulators in many scientific fields are often extremely expensive and may introduce prohibitive costs in decision making, even if they are extremely accurate. In this case, the simulator may indeed be “accurate enough,” but not under a set of realistic computational constraints. Another issue is non-localized errors that are somewhat small in absolute terms but may influence decision-making — for example, a simulator may accurately capture the relative temperature between two areas (i.e., their covariance structure), but not correctly capture their absolute magnitude, which is essential for downstream application. Finally, we note that many other OED methods rely on assumptions about the reasonableness of the prior. In contrast, MIL optimality can be interpreted as finding a design that performs well under a prior-predictive check, effectively aligning the design with the simulator-informed prior up to the limits of current modeling knowledge.
We include experiments investigating synthetic simulator error, reinforcing our theoretical analysis above, in Appendix E.
@@ 3.4 Practical Considerations: Existing Sensors & Sensor Removal
The MIL framework is rather flexible and amenable to several practical extensions to deal with existing sensors, count, or heavy-tailed data, and sensor removals.
@@
Fortunately, this issue is relatively simple to solve in our framework. The simple solution is to fix some subset of inducing points during optimization to the locations of pre-existing sensors. In this sense, we find the optimal variational approximation in which some subset of observation locations are known and fixed a priori. Note that this does not result in additional computations, nor does it interfere with solving for all $\mathbf{X}_{N}$ simultaneously. From an implementation standpoint, this can be achieved by simply not propagating the gradients with respect to existing observation locations. This process is discussed and illustrated in Appendix B.
@@
Another place where differing likelihoods may be useful is in non-Gaussian response data. For example, the sparse variational GP can also be used for other modalities, such as count data (e.g., with a Poisson likelihood) or with heavier tails (e.g., with a Student’s-$t$ likelihood). This is another rather unique aspect of our work within the Bayesian OED literature.
One practical problem common to spatio(temporal) sensor placement is that of sensor removal: suppose it is too expensive to keep all $M$ placed sensors, and we must remove $R$ of them. The MIL-optimal condition is well-suited for this, as we can simply evaluate the optimization bound with each set of $N\choose R$ sensors removed. This allows us to remove sensors with an awareness of their downstream predictive effects in non-stationary settings.
@@ 4 MIL-Optimal Sensor Placement for Spatiotemporal Processes
To tackle the spatiotemporal modeling in our problem, we will use a recent variational spatiotemporal GP, proposed by hamelijnck2021spatio. We first explain why the usual variational sparse GPs are insufficient for spatiotemporal problems (Section 4.1), before introducing sparse variational spatiotemporal GPs (Section 4.2). We then discuss the implications for computing $U_{\text{MIL}}(\cdot)$.
@@ 4.1 Initial Remarks About Spatiotemporal Processes
In many scientific applications, it is desirable for $\mathbf{X}_{N}$ to refer to a set of spatially fixed observation locations that are repeated in time, for example, to design a network of temperature sensors that observe temperature hourly. While this is a seemingly innocuous requirement, it turns out to be quite limiting for the methods discussed above.
For $U_{\text{MIL}}(\cdot)$ in particular, the usual optimization procedure becomes problematic, as the inducing points become irregularly scattered in space and time. Even after restricting to fixed spatial locations, computational problems arise, as the computation of the ELBO term in Eq. 17 scales cubically with $M$. Naïvely, for a time series of length $N_{t}$, the sparse variational GP would be implemented with $M\propto N_{t}$, leading to a potentially prohibitive $\mathcal{O}(N_{t}^{3})$ time complexity.
Thus, we formulate the following desiderata for $U_{\text{MIL}}(\cdot)$ in the spatiotemporal setting:
@@
These desiderata rule out existing approaches in multiple ways: for one, MIL-optimality is novel and requires fairly recent machinery from variational inference to optimize in a scalable fashion. Additionally, joint optimization has been, in general, difficult to achieve in BED-like approaches, leading to the widespread use of greedy algorithms like BAD or that of krause2008near. In what follows, we introduce the necessary tools to satisfy our desiderata.
@@ 4.2 Satisfying Our Desiderata
A naïve application of variational GPs to the MIL objectives is quite unlikely to succeed; only recently have new GP methods been developed that are capable of satisfying our desiderata, and the machinery remains nonstandard. Developing a computational framework for deriving MIL-optimal sensor configurations will consist of two steps: (1) spatiotemporal GP regression via stochastic differential equations (sarkka2013spatiotemporal), and (2) application of conjugate-computation variational inference (CVI) (khan2017conjugate; chang2020fast) with fixed inducing points (hamelijnck2021spatio). These two steps handle the scalability desiderata and the optimality desiderata, respectively.
@@
Overcoming the cost of GP predictive inference — which is $\mathcal{O}(N^{3})$ as computed using the kernel — is a big challenge in GP inference generally. This can be particularly problematic for long time series, where $N\propto N_{t}$ implies that datasets can grow quite quickly. For a special subset of kernels, however, it is possible to perform posterior predictive inference in linear (i.e., $\mathcal{O}(N_{t})$) time.
This approach, popularized by sarkka2013spatiotemporal, represents a GP in terms of a stochastic differential equation (SDE), where posterior predictive inference reduces to Kalman filtering and smoothing. For the moment, we will consider a temporal GP $f$, indexed by a scalar $t$, which we assume to have a stationary kernel $\kappa(t,t^{\prime})=\kappa(t-t^{\prime})$. For a wide variety of kernels, we may form an augmented state $\overline{\mathbf{f}}$ (which often consists of the value of $f$ and some time derivatives of $f$) and a corresponding linear time-invariant SDE that describes the GP $f$. This SDE is usually marginalized to a discrete-time linear dynamical system,
@@
where $\boldsymbol{\mathbf{q}}_{n}\overset{\mathrm{iid}}{\sim}\mathcal{N}(\mathbf{0},\boldsymbol{\mathbf{Q}}_{n})$ and $\varepsilon_{t_{n}}\overset{\mathrm{iid}}{\sim}\mathcal{N}(0,\sigma_{n}^{2})$. The parameters of this linear dynamical system (i.e., the matrices $\mathbf{A}_{n}$ and $\mathbf{H}$, and the noise (co)variances $\mathbf{Q}_{n}$ and $\sigma_{n}^{2}$) are determined directly by the choice of kernel $\kappa(t-t^{\prime})$.
GPs that can be written using Eq. 26 are referred to as Markovian GPs, or sometimes state-space GPs. The latter term may be confused with state-space models in which GPs describe the state transition; thus, we will avoid it. Although not all GPs are Markovian, the state-space representation can be derived exactly for many standard choices of kernels – for example, kernels from the Matérn family – and may be done approximately for many others (sarkka2013spatiotemporal; sarkka2019applied).
After converting to this representation, exact linear time inference can proceed via the Kalman filtering and smoothing algorithms (sarkka2023bayesian). Furthermore, the marginal likelihood can be computed in the filtering step, so the $\mathcal{O}(N_{t})$ computation can also be used in optimization. In the spatiotemporal case, we can use infinite-dimensional filtering techniques, which also scale linearly with $N_{t}$ (sarkka2013spatiotemporal).
Alas, the use of Markovian GPs alone does not solve the rest of our desiderata. In particular, incorporating variational objectives that solve MIL-optimality is not straightforward. Fortunately, a recent application of CVI has shown this to be possible.
@@
Despite their impressive $\mathcal{O}(N_{t})$ scaling, spatiotemporal Markovian GPs still struggle with $\mathcal{O}(N_{s}^{3})$ inference, where $N_{s}$ is the number of spatial locations. This can be understood by recognizing that the state dimension scales linearly with $N_{s}$, and Kalman filtering/smoothing algorithms have cubic time complexity in their state dimension. This and non-conjugate likelihoods have led to several works on sparse and variational approaches to Markovian GPs (wilkinson2021sparse; adam2020doubly; chang2020fast).
The main additional assumption we must make is that the GP kernel is separable, i.e.,
@@
Probabilistically, this leads to a factorization across time, and computationally, it leads to a state space with a Kronecker-like structure.
Using this assumption, hamelijnck2021spatio apply CVI to obtain a spatiotemporal sparse variational GP (ST-SVGP) with some remarkable properties. CVI reformulates variational inference as approximate Bayesian inference with respect to a conjugate exponential family model and allows efficient natural-gradient updates of the variational parameters. In the case of ST-SVGPs, we fix the inducing point locations to a $M_{s}\times N_{t}$ spatiotemporal grid. The CVI algorithm then proceeds efficiently in linear-time along the temporal dimension (via Kalman filtering) using the $M_{s}N_{t}$ pseudo observations, followed by natural gradient descent on the pseudo-likelihood parameters. Remarkably, the resulting predictive distribution is exactly equivalent to that of a traditional VGP with inducing points on a spatiotemporal grid. With $\boldsymbol{\mu}_{t}$ and $\mathbf{A}_{t}$ denoting the inducing point distributions at time $t$, Eqs. 18 and 19 together with Eq. 28 then imply
@@
where
@@
$\boldsymbol{\Sigma}_{t}$ $=\mathbf{K}_{NM}\mathbf{K}_{MM}^{-1}\mathbf{A}_{t}\mathbf{K}_{MM}^{-1}\mathbf{K}_{MN}+\mathbf{K}_{NN}.$ (31)
Thus, we obtain a predictive formula depending on $\mathbf{m}_{t}$ and $\mathbf{A}_{t}$.
What remains, then, is obtaining $\mathbf{m}_{t}$ and $\mathbf{A}_{t}$ from the test data. It is clear that the value of $\mathbf{m}_{t}$ on training data will typically be different from its value on testing data. However, it is more realistic to assume that the covariance structure $\mathbf{A}_{t}$ remains intact. For all of our experiments, we reuse $\mathbf{A}_{t}$ directly and obtain $\mathbf{m}_{t}$ through a single pass of CVI on the test data, keeping $\mathbf{A}_{t}$ fixed. Notably, different choices in determining $\mathbf{A}_{t}$ exist; for example, one may use additional simulation data that directly emulate the observed data.

### v2

@@ 3.2 Connection to Sparse Variational Gaussian Processes
Our first assumption regards signal-to-noise ratios, allowing us to replace the inducing point means $\boldsymbol{\mathbf{\mu}}=\mathbb{E}[q(\boldsymbol{\mathbf{u}})]$ with observations. Recall the analytic solution of $\boldsymbol{\mathbf{\mu}}$,
$\boldsymbol{\mathbf{\mu}}=\sigma^{-2}\boldsymbol{\mathbf{K}}_{MM}\mathopen{}\mathclose{{\left(\boldsymbol{\mathbf{K}}_{MM}+\sigma^{-2}\boldsymbol{\mathbf{K}}_{MN}\boldsymbol{\mathbf{K}}_{NM}}}\right)^{-1}\boldsymbol{\mathbf{K}}_{MN}\boldsymbol{\mathbf{y}},$ (21)
where $\sigma^{2}$ is determined as the estimated noise variance plus a regularizer $\operatorname{tr}{\boldsymbol{\mathbf{K}}_{\text{post}}}/N$. So long as the sparse variational approximation is good, we expect $\operatorname{tr}{\boldsymbol{\mathbf{K}}_{\text{post}}}/N$ to be small, and thus $\sigma^{2}$ to be close to the estimated noise level. Then, with a high signal-to-noise ratio in the simulation, an observation $y(\boldsymbol{\mathbf{z}}_{m})$ will approximately be $\mu_{m}$.
If we additionally assume a high signal-to-noise ratio in observations of the real process, we can thus approximate the corresponding ${\mu}_{m}$. This is a surprisingly realistic assumption in many settings: for example, in temperature observations, noise may be on the scale of a few tenths of a degree Celsius.
@@
We proceed with a perturbative analysis, investigating the effect of data perturbation $\delta_{\boldsymbol{\mathbf{Y}}}\triangleq\tilde{\boldsymbol{\mathbf{Y}}}_{N}-\boldsymbol{\mathbf{Y}}_{N}.$ In the collapsed bound Eq. 17, the data enter only through the data fit term
$D(\mathbf{Y}_{N})\triangleq\frac{1}{2}\boldsymbol{\mathbf{Y}}_{N}^{\top}(\boldsymbol{\mathbf{Q}}_{NN}+\sigma^{2}_{\text{obs}}\mathbf{I})^{-1}\boldsymbol{\mathbf{Y}}_{N}.$ (22)
@@
through perturbations of the data fit term. Since $\mathcal{L}_{V}$ depends on $\mathbf{Y}_{N}$ only through $D(\mathbf{Y}_{N})$, we have $\Delta\mathcal{L}=D(\mathbf{Y}_{N}+\delta_{\mathbf{Y}})-D(\mathbf{Y}_{N})$; since $D(\cdot)$ is a quadratic form,
$\Delta\mathcal{L}$ $=\mathopen{}\mathclose{{\left[\nabla D(\boldsymbol{\mathbf{Y}}_{N})}}\right]^{\top}\delta_{\boldsymbol{\mathbf{Y}}}+D(\delta_{\boldsymbol{\mathbf{Y}}})$ (24)
where $\nabla D(\boldsymbol{\mathbf{Y}}_{N})=(\boldsymbol{\mathbf{Q}}_{NN}+\sigma^{2}_{\text{obs}}\mathbf{I})^{-1}\boldsymbol{\mathbf{Y}}_{N}$.
@@
    A linear term, $\mathopen{}\mathclose{{\left[\nabla D(\boldsymbol{\mathbf{Y}}_{N})}}\right]^{\top}\delta_{\boldsymbol{\mathbf{Y}}}$. There are two notable ways for this term to be small. The first is that both $\nabla D(\boldsymbol{\mathbf{Y}}_{N})$ and $\delta_{\boldsymbol{\mathbf{Y}}}$ are small, which can occur when the simulator is accurate (hence $\delta_{\boldsymbol{\mathbf{Y}}}$ is small). Potentially more interesting is the case in which $\nabla D(\boldsymbol{\mathbf{Y}}_{N})$ and $\delta_{\boldsymbol{\mathbf{Y}}}$ are nearly orthogonal under the metric induced by $(\mathbf{Q}_{NN}+\sigma^{2}_{\text{obs}}\mathbf{I})^{-1}$. This situation may be interpreted as the real data deviating from the simulator in directions that are well captured by the covariance structure.
@@
    A quadratic term, $D(\delta_{\boldsymbol{\mathbf{Y}}})$. This term is small precisely when $\delta_{\boldsymbol{\mathbf{Y}}}$ has small energy under the inverse covariance $(\mathbf{Q}_{NN}+\sigma^{2}_{\text{obs}}\mathbf{I})^{-1}$, that is, when the perturbation lies primarily in directions of high posterior uncertainty.
In total, this indicates that the simulator error is structured in such a way that it it primarily lies in directions of high variance under the covariance $(\boldsymbol{\mathbf{Q}}_{NN}+\sigma^{2}_{\text{obs}}\mathbf{I})$. Equivalently, the error has low energy under the corresponding inverse covariance. Unfortunately, it is difficult to test this property directly, but it can be informed via domain expertise or empirical evaluations. This limitation is inherent to our “pre-observation” setting, as no reliable statistical test can be designed without data learned from subsequent deployment. Nonetheless, we provide the following heuristics to help understand when this assumption is plausible:
@@
    High signal-to-noise ratio: The outputs of the simulator should dominate internal numerical noise or model uncertainty. This property can be checked empirically by computing the signal-to-noise ratios across simulated realizations or by estimating noise levels during GP training.
@@
    Non-localized errors: The simulator may exhibit errors, but those should be spatially diffuse rather than concentrated in specific regions. Our perturbative analysis (Section 3.2) shows that, under small perturbations, non-localized errors tend to have little effect on the ELBO and the resulting inducing points.
@@
    Coherence with sparse observations: If a limited number of real-world measurements are available, comparing them to simulator outputs can help establish empirical plausibility. In particular, modest correlation or bounded deviations support the assumption that simulation-informed placement remains meaningful.
Note that small and non-localized errors may not be sufficient for a simulator to be useful in real-world decision making without the incorporation of real data. One potential reason for this is computational constraints: simulators in many scientific fields are often extremely expensive and their direct use may introduce prohibitive costs in decision time, even if they are highly accurate. In this case, the simulator may indeed be “accurate enough,” but not under realistic computational constraints.
Another limitation arises when non-localized errors that are small in absolute terms but nonetheless influence decision-making. For example, a simulator may accurately capture the relative temperature between two areas (i.e., their covariance structure), yet fail to capture their absolute magnitudes, which is essential for downstream applications.
Finally, we note that many other OED methods rely on assumptions about the reasonableness of the prior. In contrast, MIL optimality can be interpreted as finding a design that performs well under a prior-predictive check, effectively aligning the design with the simulator-informed prior up to the limits of current modeling knowledge.
We include experiments investigating synthetic simulator error, reinforcing our theoretical analysis above, in Appendix E (Waxman et al., 2025).
@@ 3.4 Practical Considerations: Existing Sensors & Sensor Removal
The MIL framework is flexible and amenable to several practical extensions to deal with existing sensors, count, or heavy-tailed data, and sensor removals.
@@
Fortunately, this issue is relatively simple to solve within our framework. The simple solution is to fix some subset of inducing points during optimization to the locations of pre-existing sensors. In this sense, we find the optimal variational approximation in which some subset of observation locations is known and fixed a priori. Note that this does not result in additional computations, nor does it interfere with solving for all $\mathbf{X}_{N}$ simultaneously. From an implementation standpoint, this can be achieved by simply not propagating the gradients with respect to existing observation locations. This process is discussed and illustrated in Appendix B (Waxman et al., 2025).
@@
Differing likelihoods may also be useful for non-Gaussian response data. For example, the sparse variational GP can also be used for other modalities, such as count data (e.g., with a Poisson likelihood) or with heavier tails (e.g., with a Student’s-$t$ likelihood). This is another unique aspect of our work within the Bayesian OED literature.
One practical problem common to spatio(temporal) sensor placement is that of sensor removal: suppose it is too expensive to keep all $N$ placed sensors, and we must remove $R$ of them. The MIL-optimal condition is well-suited for this, as we can simply evaluate the optimization bound with each set of $N\choose R$ sensors removed, allowing us to remove sensors with an awareness of their downstream predictive effects.
@@ 4 MIL-Optimal Sensor Placement for Spatiotemporal Processes
To tackle the spatiotemporal modeling requirements of our problem, we will use a recent variational spatiotemporal GP model proposed by Hamelijnck et al. (2021). We first explain why the usual variational sparse GPs are insufficient for spatiotemporal settings (Section 4.1), before introducing sparse variational spatiotemporal GPs (Section 4.2). We then discuss the implications for computing $U_{\text{MIL}}(\cdot)$.
@@ 4.1 Initial Remarks About Spatiotemporal Processes
In many scientific applications, it is desirable for $\mathbf{X}_{N}$ to refer to a set of spatially fixed observation locations, repeatedly observed over time, for example, when designing a network of temperature sensors that record measurements hourly. While this is a seemingly innocuous requirement, it turns out to be quite limiting for existing methods.
For $U_{\text{MIL}}(\cdot)$ in particular, the usual optimization procedure becomes problematic, as the inducing points become irregularly distributed in space and time. Even after restricting to fixed spatial locations, computational problems arise, as the computation of the ELBO term in Eq. 17 scales cubically with the number of inducing points $M$. Naïvely, for a time series of length $N_{t}$, the sparse variational GP would be implemented with $M\propto N_{t}$, leading to potentially prohibitive $\mathcal{O}(N_{t}^{3})$ computational complexity.
We thus formulate the following desiderata for $U_{\text{MIL}}(\cdot)$ in spatiotemporal settings:
@@
These desiderata rule out existing approaches in multiple ways: for one, MIL-optimality is novel and requires fairly recent machinery from variational inference to optimize in a scalable fashion. Additionally, joint optimization has been, in general, difficult to achieve in BED-like approaches, leading to the widespread use of greedy algorithms like BAD or that of Krause et al. (2008). In what follows, we introduce the necessary tools to satisfy our desiderata.
@@ 4.2 Satisfying Our Desiderata
A naïve application of variational GPs to the MIL objectives is quite unlikely to succeed; only recently have new GP methods been developed that can satisfy our desiderata, and the machinery remains nonstandard. Developing a computational framework for deriving MIL-optimal sensor configurations will consist of two steps: (1) spatiotemporal GP regression via stochastic differential equations (Särkkä et al., 2013), and (2) application of conjugate-computation variational inference (CVI) (Khan and Lin, 2017) with fixed inducing points (Hamelijnck et al., 2021). These two steps handle the scalability desiderata and the optimality desiderata, respectively.
@@
Overcoming the cost of GP posterior inference — which is $\mathcal{O}(N^{3})$ as computed using the kernel — is a big challenge in GP inference generally. This can be particularly problematic for long time series, where $N\propto N_{t}$ implies that datasets can grow quite quickly. For a special subset of kernels, however, it is possible to perform posterior predictive inference in linear (i.e., $\mathcal{O}(N_{t})$) time.
This approach, popularized by Särkkä et al. (2013), represents a GP in terms of a stochastic differential equation (SDE), where posterior inference reduces to Kalman filtering and smoothing. For the moment, we will consider a temporal GP $f$, indexed by a scalar $t$, which we assume to have a stationary kernel $\kappa(t,t^{\prime})=\kappa(t-t^{\prime})$. For many common kernels, we may form an augmented state $\overline{\mathbf{f}}$ (which often consists of the value of $f$ and some time derivatives of $f$) and a corresponding linear time-invariant SDE that describes the GP $f$. This SDE is usually marginalized to a discrete-time linear dynamical system, with linear observations,
@@
where $\boldsymbol{\mathbf{q}}_{n}\overset{\mathrm{iid}}{\sim}\mathcal{N}(\mathbf{0},\boldsymbol{\mathbf{Q}}_{n})$ and $\varepsilon_{t_{n}}\overset{\mathrm{iid}}{\sim}\mathcal{N}(0,\sigma_{n}^{2})$. The parameters of this linear dynamical system (i.e., the matrices $\mathbf{A}_{n}$ and $\mathbf{H}$, and the noise (co)variances $\mathbf{Q}_{n}$ and $\sigma_{n}^{2}$) are determined directly by the choice of kernel $\kappa(t-t^{\prime})$ and the discretization interval.
GPs that can be written using Eq. 25 are referred to as Markovian GPs, or sometimes state-space GPs. The latter term may be confused with state-space models in which GPs describe the state transition; thus, we will avoid it. Although not all GPs are Markovian, the state-space representation can be derived exactly for many standard choices of kernels – for example, kernels from the Matérn family – and may be approximated for many others (Särkkä et al., 2013).
After converting to this representation, exact linear time inference can proceed via Kalman filtering and smoothing algorithms (Särkkä and Svensson, 2023). Furthermore, the marginal likelihood can be computed in the filtering step, so that the evaluation of likelihood-based optimization objectives also scales as $\mathcal{O}(N_{t})$. In the spatiotemporal case, we can use infinite-dimensional filtering techniques, which also scale linearly with $N_{t}$ (Särkkä et al., 2013).
However, the use of Markovian GPs alone does not solve the rest of our desiderata. In particular, incorporating variational objectives that solve MIL-optimality is not straightforward within this framework. Fortunately, a recent application of CVI has shown this to be possible.
@@
Despite their impressive $\mathcal{O}(N_{t})$ scaling, spatiotemporal Markovian GPs still incur an $\mathcal{O}(N_{s}^{3})$ computational cost, where $N_{s}$ is the number of spatial locations. This can be understood by recognizing that the state dimension scales linearly with $N_{s}$, and Kalman filtering/smoothing algorithms have cubic time complexity in their state dimension. This limitation and the presence of non-conjugate likelihoods have led to several works on sparse and variational approaches to Markovian GPs (Wilkinson et al., 2021; Chang et al., 2020).
The main additional assumption we must make is that the GP kernel is separable:
@@
Probabilistically, this induces a tensor-product structure in the spatiotemporal covariance, while computationally it leads to state-space models Kronecker-like structure.
Using this assumption, Hamelijnck et al. (2021) apply CVI to obtain a spatiotemporal sparse variational GP (ST-SVGP) with favorable computational properties. CVI reformulates variational inference as approximate Bayesian inference with respect to a conjugate exponential family model and allows efficient natural-gradient updates of the variational parameters. In the case of ST-SVGPs, we fix the inducing point locations to a $M_{s}\times N_{t}$ spatiotemporal grid. The CVI algorithm then proceeds efficiently in linear-time along the temporal dimension (via Kalman filtering), followed by natural gradient updates of the pseudo-likelihood parameters. Remarkably, the resulting predictive distribution is exactly equivalent to that of a traditional VGP with inducing points on a spatiotemporal grid.
With $\boldsymbol{\mu}_{t}$ and $\mathbf{A}_{t}$ denoting the variational mean and covariance of the inducing variables at time $t$, Eqs. 18 and 19 together with Eq. 27 then imply
@@
where writing $\mathbf{K}_{NN}$, $\mathbf{K}_{NM}$, and $\mathbf{K}_{MM}$ for the appropriate kernel blocks at time $t$,
@@
$\boldsymbol{\Sigma}_{t}$ $=\mathbf{K}_{NN}-\mathbf{K}_{NM}\mathbf{K}_{MM}^{-1}\mathbf{K}_{MN}+\mathbf{K}_{NM}\mathbf{K}_{MM}^{-1}\mathbf{A}_{t}\mathbf{K}_{MM}^{-1}\mathbf{K}_{MN}.$ (30)
What remains, then, is obtaining $\mathbf{m}_{t}$ and $\mathbf{A}_{t}$ from the test data. It is clear that the value of $\mathbf{m}_{t}$ on the training data will typically be different from its value on the testing data. However, it is more realistic to assume that the covariance structure $\mathbf{A}_{t}$ remains intact. For all of our experiments, we reuse $\mathbf{A}_{t}$ directly and obtain $\mathbf{m}_{t}$ through a single pass of CVI on the test data, keeping $\mathbf{A}_{t}$ fixed. Notably, different choices for determining $\mathbf{A}_{t}$ exist; for example, one may use additional simulation data that directly emulate the observed data.